If the difference between the measures of the two smaller angles of a right triangle is 8 degrees, what is the measure of the smallest angle?

Answer by DrBeeee(684)

Let x = one of the angles
Let y = the other angle
We have two equations
(1) x + y = 90, and
(2) x - y = 8
To solve just add (1) and (2) and get
(3) 2*x + 0 = 98 or
(4) x = 49
Then from (1) we have
(5) 49 + y = 90 or
(6) y = 90 - 49 or
(7) y = 41
Let's check with (2).
Is (49 - 41 = 8)?
Is (8 = 8)? Yes
Answer: The smallest angle is 41 degrees.
REM: You need two equations to find two unknowns.